OpenStudy (anonymous):

pipe A can fill a tank in 4 hours. pipe B can fill in 9 hours less then the time it takes pipe C, a drain pipe, to empty the tank. When all 3 pipes are open, it takes 2 hours to fill the tank. How much time is required for pipe C to empty the tank if pipes A and B are closed. CAN SOMEONE SHOW ME THE STEPS FOR IT?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

Given Pipe A fills tank in 4 hours Pipe C empties tank in C hours Pipe B fills tank in C-9 hours -- this means that every hour Pipe A filles 1/4 of tank Pipe C empties 1/c of tank Pipe B fills 1/(c-9) of tank When all pipes are open (ie working together) they fill the tank in 2 hours soo..... "every hour", working together 1/2 the tank is filled. Pipe A + pipe B -Pipe C = \[C=\frac{1}{4}+\frac{1}{c-9}-\frac{1}{c}=\frac{1}{2} \]\[\frac{1}{c-9}-\frac{1}{c}=\frac{1}{2}-\frac{1}{4}=\frac{1}{4}\] Subtract 1/4 from both side

OpenStudy (anonymous):

Thank You very much. So what would be the final answer?

OpenStudy (anonymous):

\[\frac{c}{c*(c-9}-\frac{c-9}{(c-9)*c}=\frac{1}{4}\]get common denominator\[\frac{c-(c-9)}{c*(c-9)}=\frac{1}{4}\]\[\frac{9}{c ^{2}-(9 \times c)}=\frac{1}{4}\]combine and distribute terms\[36=c ^{2}-9c\]\[0=c ^{2}-(9 \times c) -36\]

OpenStudy (anonymous):

factor 0=(c-12)*(c+3) so c=12, or c=-3, since solution must be positive c=-3 is extraneous answer: c=12 -- You could use use quadratic solution, a=1,b=-9, c=-36 \[ c=\frac{\sqrt{(-9)^{2}-(4\times 1\times(-36)}}{2 \times 1}\]or \[c=\frac{9-\sqrt{(-9)^{2}-(4 \times 1 \times(-36) }}{2 \times 1}\]

OpenStudy (anonymous):

sorry my computer go so slow an I can answer u either

OpenStudy (anonymous):

oh wow ur a God. its all good thank u very much for helping me out. :)

OpenStudy (anonymous):

when I been working ,if I answer you , I will lost every thing and I have restart over again

OpenStudy (anonymous):

any more help?

OpenStudy (anonymous):

1 more question?

OpenStudy (anonymous):

I will try

OpenStudy (anonymous):

Together Maria and Johannes can pick a bushel of apples in 16 min. Maria takes 4 min. more to pick a bushel than johannes does. Find the time it takes each person to pick a bushel of apples alone?

OpenStudy (anonymous):

give me time to working on it , and renember I anable your question when I been typeing

OpenStudy (anonymous):

its cool thank you very much for doing this... its my first time and im stoked that i can find help!

OpenStudy (anonymous):

can we go anothe page can ca get more medal?

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

u don't have to write all just post hi and i know how answer

OpenStudy (anonymous):

let go